OpenStudy (anonymous):

log (x^2-36) -1= log (x+6) Solve for X.

OpenStudy (anonymous):

Okay i know that you have to do 10^log (x^2-36) -1= 10^log (x+6)

OpenStudy (anonymous):

and then the logs cancel out,

OpenStudy (anonymous):

The base is 10, so we don't need to take that out right now. First put the -1 on the other side. log(x^2-36)-1=log(x+6) log(x^2-36)-log(x+6)=1 log[(x^2-36)/(x+6)]=1 Heres a clearer view.  Do you understand so far?